OpenStudy (anonymous):

simplify the expression csc(-x) 1+tan^2(x)

myininaya (myininaya):

Hint: for the first one. sin(x) is an odd function which means sin(-x)=-sin(x)

myininaya (myininaya):

Hint: for the second one. Think about your Pythagorean identities.

OpenStudy (anonymous):

So is it sin(x)tan(x)?

myininaya (myininaya):

I'm not sure how you got csc(-x) is sin(x)tan(x)

myininaya (myininaya):

recall \[\csc(x)=\frac{1}{\sin(x)}\]

OpenStudy (anonymous):

I have no idea how to do this. lol

myininaya (myininaya):

\[\csc(-x)=\frac{1}{\sin(-x)}=\frac{1}{?}\] can you use anything I have mentioned here to rewrite sin(-x)?

OpenStudy (anonymous):

The equation is over each other like csc(-x) over 1+tan^2(x) and the choices given are tan(x) -cos(x)tan(x) -cos(x)cot(x) sin(X)tan(x)

myininaya (myininaya):

ok so you meant to write a fraction

OpenStudy (anonymous):

yes!

myininaya (myininaya):

anyways the expression I was asking about can you rewrite it

myininaya (myininaya):

i dare you to use the hint I gave above

myininaya (myininaya):

1/sin(-x) =?

OpenStudy (anonymous):

1 over sin(x) ?

myininaya (myininaya):

sin(-x) doesn't equal sin(x)

myininaya (myininaya):

sin(-x) equals -sin(x) since sin(x) is an odd function

myininaya (myininaya):

this was the first hint I mentioned

myininaya (myininaya):

\[\csc(-x)=\frac{1}{\sin(-x)}=\frac{1}{-\sin(x)}\]

myininaya (myininaya):

now can you use your Pythagorean identities to rewrite the bottom of your fraction as one term that is I'm asking you to rewrite 1+tan^2(x)

OpenStudy (anonymous):

is that sec^2(X)?

myininaya (myininaya):

yes

myininaya (myininaya):

so you have this so far: \[\frac{\csc(-x)}{1+\tan^2(x)} \\ =\csc(-x) \cdot \frac{1}{1+\tan^2(x)} \\ =\frac{1}{\sin(-x)} \cdot \frac{1}{1 +\tan^2(x)} \\ = \frac{1}{-\sin(x)} \cdot \frac{1}{\sec^2(x)}\]

myininaya (myininaya):

now recall that sec(x) and cos(x) are reciprocal functions of one another

myininaya (myininaya):

that is \[\frac{1}{\sec(x)}=\cos(x)\]

myininaya (myininaya):

try that to use that here

OpenStudy (anonymous):

So -cos(x)tan(x)?

myininaya (myininaya):

lets take little steps

myininaya (myininaya):

1/sec^2(x)=?

OpenStudy (anonymous):

cos^2(x)?

myininaya (myininaya):

yes so we have \[\frac{1}{-\sin(x)} \cdot \cos^2(x) \\ \frac{1}{-\sin(x) } \cos(x) \cos(x) \] now rearrange some things so you can see this is equivalent to one of your choices

myininaya (myininaya):

\[- \frac{\cos(x)}{\sin(x)} \cos(x)\]